# Structural Chirality and Natural Optical Activity across the -to‑ Phase Transition in SiO2 and AlPO4 from first-principles

### v1

@@ Abstract
Natural optical activity (NOA), the ability of a material to rotate the plane of polarized light, has traditionally been associated with structural chirality. However, this relationship has often been oversimplified, leading to conceptual misunderstandings, particularly when attempts are made to directly correlate structural handedness with optical rotatory power. In reality, the relationship between chirality and NOA is more nuanced: optical activity can arise in both chiral and achiral crystal structures, and the sign of the rotation cannot necessarily be inferred from the handedness of the space group. In this work, we conduct a first-principles investigation of natural optical activity in SiO2 and AlPO4 crystals, focusing on their enantiomorphic structural phase transition from high-symmetry hexagonal ( or ) to low-symmetry trigonal ( or ) space groups. This transition, driven by the condensation of a zone-center phonon mode, reverses the screw axis type given by the space group symbol while leaving the sign of the optical activity unchanged. By following the evolution of the structure and the optical response along the transition pathway, we clarify the microscopic origin of this behavior. We demonstrate that the sense of optical rotation is determined not by the nominal helicity of the screw axis given in the space group symbol, but by the atomic-scale helicity of the most polarizable atoms of the structure.
@@ I Introduction
While it is a well-known fact that NOA is frequently linked with structural chirality, it is currently understood that chirality is a sufficient, albeit not necessary, condition for NOA to occur, given that achiral materials can also display NOA characteristics. [17, 18, 13]. This historical association led to the widespread belief that the chirality of a material’s space group directly determines the handedness of its optical rotation [19, 20]. Although this correlation may hold in many cases, it is more nuanced: the direction of optical rotation does not necessarily follow the handedness/helicity given in the space group symbol. [21, 22]
Measurements of optical activity are a common practice for organic chiral molecules (e.g., sugars or tartaric acid) and, although comparatively weaker, remain measurable in inorganic crystals such as low-quartz and berlinite [21]. Quartz-crystal (SiO2) adopts a trigonal structure (enantiomorphic space groups, 21 or 21) at room temperature (“-quartz”) and transforms sharply to hexagonal “-quartz” (enantiomorphic space groups, or respectively) at 573 °C; both polymorphs are chiral despite being built from achiral SiO4 tetrahedra. Berlinite (AlPO4), the structural analogue of quartz, similarly undergoes an transition at approximately 586 °C, passing from trigonal () symmetry to hexagonal () without disrupting the network of corner-sharing tetrahedra [23, 24, 25]. In such transitions, the sense of the principal screw axis of the space group symbol reverses sign; however, the sense of the optical activity of the crystal remains unchanged [26, 22]. Initially, it might appear counterintuitive, but the direction of rotation for the helix composed of the most polarizable atoms (which is unchanged throughout the transition) actually determines the direction of optical activity, as shown by Glazer and Stadnicka. [21]. This explains why the optical activity is preserved despite the reversal of the principal screw axis in these crystal structures.
@@ III.1 Structural transition of SiO2 and AlPO4
Recently, an algorithm has been proposed to identify chiral displacive transitions from achiral to chiral space groups [27], following the works of Fava et al. [36] for K3NiO2. Applying this workflow [27] to SiO2 and AlPO4, we find that a displacive phase transition from an achiral to a chiral space group is not allowed by pseudosymmetry. In other words, this means that an achiral phase of SiO2 or AlPO4 cannot be built from their enantiomorphic crystal structure if one wants a continuous small-atom distortion pattern connecting the two.
@@
However, this pseudosymmetry search predicts that the high symmetry right-handed screw-axis structure can be linked to the low-symmetry 21 left-handed screw-axis structure [24, 25] through the irrep of the zone center [37, 28] (or similarly between the and the space groups). This means that those two space groups can form a low-symmetry/high-symmetry pairs for SiO2 or AlPO4. This kind of spontaneous symmetry breaking is possible since a chiral irreducible representation (or order parameter) can be formally defined [38], so that a group-subgroup relationship between two phases with distinct chirality is physically possible [39].
@@
From our DFT calculations, we confirm the presence of an unstable phonon mode at the point in the phase of SiO2, as referenced by [40]. By checking the corresponding phonon eigenvector symmetry, we found that it has the irrep, i.e. in agreement with the symmetry analysis which links the right-handed screw of the high-symmetry phase and the left-handed screw of the low-symmetry phase. Such a mode is neither Raman nor infrared active, i.e. it is silent as reported some times ago through hyper-Raman [41] or inelastic neutron scattering [42]. Hence, these results go in the direction of the displacive character of the transition [43], helping clarify previous discussions on its nature [44]. The calculated structural data of the converged structure computed with a -mesh of 666 can be found in Table 1. From our DFPT calculations, we found that the imaginary frequency of the mode has an amplitude of cm-1. The condensation and relaxation of the eigenvector gives an energy difference between the high and low temperature phases of meV per 9 atoms within the unit cell.
| Atom | x | y | z | |||
| Si (3a) | 0.0000 | 0.5000 | 0.3333 | 0.0000 | 0.0000 | |
| O (6c) | 0.2080 | 0.4160 | 0.8333 | 0.0606 |
@@
A symmetry adapted mode analysis using the isodistort software[45, 46] of our relaxed low-symmetry phase reveals that the distortion decomposes into and irreps where the amplitude of the mode is marginal (0.03 Å) with respect to the leading mode (1.00 Å) which drives the transition. The mode induces in-plane displacements of the Si atoms and both out-of-plane and in-plane displacements of the O atoms. The displacements are such that the SiO4 tetrahedra rotate as fairly rigid units (and do not distort significantly), as illustrated in Figure 1(c). As shown in Fig. 1(a-b), the in-plane projection of the displacements induces a coordinated counter-clockwise and clockwise movement between the neighboring helical centers. This movement breaks the -fold screw axis at the corners of the unit cells that transform into centers while preserving the -fold screw axis at the interior of the unit cell.
@@
Moving now to the case of AlPO4 crystal and when studying its 22 phase, we can observe a similar crystal structure to that of the 22 of SiO2 [47, 48] with the exception that the -axis is essentially doubled due to the ordering of Al + P over the atom sites (as discussed below). The crystal structure is shown in Fig. 2. The calculated structural data of the converged AlPO4 22 structure can be found in Table 2. Similarly to the previous SiO2 case, we also obtain an unstable phonon mode at the point of the right-handed high-symmetry phase with an amplitude of cm-1 and the irrep [48]. This mode, hence, gives the link between the right-handed high-symmetry and the left-handed low-symmetry chiral phases, again similarly to the case of SiO2 but with doubled cell along the z direction. The condensation and relaxation of the eigenvector of this unstable mode gives, in this case, an energy gain of meV per 18 atoms unit cell, which is comparable to SiO2 when normalized to the number of atoms. The relaxed structure we obtain is fully consistent with previous works [49, 47].
Similar conclusions to those drawn in the previous example can be reached by analyzing the distortion induced by the mode, which causes in-plane displacements of the Al and P atoms and both in-plane and out-of-plane displacements of the O atoms as in the case of the quartz crystal. The symmetry adapted mode decomposition gives amplitudes of 0.08 Å for the irrep and 1.57 Å for the irrep.
@@ III.2 Evolution of the natural optical activity
The results are summarized in Fig. 3 and Table 3. The first interesting remark emerging from our calculations is the opposite sign of the NOA in the high-symmetry high-temperature phases of SiO2 and AlPO4, despite sharing the same enantiomorphic space group. Our first-principles calculations predict a positive rotatory power for the and phases of SiO2 for light traveling along its optical axis, in agreement with Ref. [11] (despite the sign of the screw axis given in this space group symbol) and a negative rotatory power for light traveling along its optical axis for the and phases of AlPO4. The numerical values we obtained for SiO2 are in good agreement with the values encountered in the literature [50] and the change of rotatory sign between SiO2 and AlPO4 is also consistent with a previous report by Glazer and Stadnicka [21]. This finding underscores a crucial concept: the sign of a crystal’s natural optical activity links to its structural chirality, though it might not align with the helical handedness indicated by the space group symbol [21, 51].
@@
We can trace and explain the difference in rotatory power by examining the crystal structures of both materials. As mentioned in the previous section and now shown in Fig. 4, the atomic positions of AlPO4 in its setting are very similar to those of the SiO2 crystal in its setting. In fact, when we perform a computational experiment by substituting the Si atoms into the Al and P sites of the AlPO4 structure, and calculate the optical activity, we obtain a rotatory power with the opposite sign to that of SiO2 in its phase and that closely resembles that found for AlPO4 in its phase. Furthermore, upon relaxing the atomic positions in this modified structure, the system naturally evolves into the phase of SiO2 although initially a symmetry is detected by Abinit. Although the one-to-one alternation of Al and P atoms and the non-equivalence of the O atoms induces a right-handed screw axis of the material as discussed in Ref. [21, 47], Fig. 4(d) shows that a non-symmetric left-handed helix similar to that observed in 22 SiO2 is indeed present in the crystal. Although not perfectly symmetric, this helix of the most polarizable oxygen sites dominates the optical activity of these phases explaining its rotatory power [21, 51]. Moreover, the observation that substituting Si atoms at the Al and P sites does not significantly alter the rotatory power suggests that these atomic species contribute little to the optical activity, which mainly arises from the more polarizable oxygen atoms.
The second interesting observation is that the condensation of the chiral distortion decreases the value of the rotatory power in these types of transitions where the screw-axis type is reverted. Indeed, as shown in Fig. 3 we observe a linear decrease of the rotatory power as we condense the distortion. This can be explained by looking at the distortions shown in Fig. 1 and Fig. 2 that tend to disrupt the helical arrangement of the O atoms. One might expect that further increasing the amplitude of the distortion could eventually revert the sign of the optical activity; however, our calculations show that, for instance, an amplitude 50% larger than the optimal value, recovers rotatory power values comparable to those of the undistorted high-symmetry phase. Moreover, this increase in distortion amplitude is accompanied by a significant increase in the Landau energy as shown in Fig. 3 (b), indicating that such configurations are physically inaccessible.
@@ IV Conclusion
In this work, we performed an ab-initio analysis of the structural transitions and associated changes in the natural optical activity between the high-symmetry () and low-symmetry () phases of SiO2 and AlPO4. Our first-principles calculations show that both materials undergo a displacive phase transition driven by the softening of a phonon mode at the zone center with symmetry. This mode induces a chiral distortion that reverses the sign of the principal screw-axis of the crystal high-symmetry space group.
Despite this reversal, the sense of the natural optical activity of the materials is preserved throughout the transition. AlPO4 and SiO2 therefore provide a useful illustration of the principle that it is the helicity of the most polarizable atoms in a structure that determines the sign of its natural optical activity [21, 51]. Upon cooling through its phase transition, the screw-sense of the imperfectly symmetric helix of oxygen sites is not changed (although this is not obvious from the space group symmetries of the and phases where the helix handedness given by the space group symbol is reversed), preserving therefore the optical activity of the system.
@@
Furthermore, our analysis of the chiral distortion pathway indicates a linear decrease in the magnitude of the rotatory power as the mode progresses from the high-symmetry to the low-symmetry phase. This trend suggests that the symmetry-lowering introduced by the distortion reduces the helicity of the structure (as shown in Figure 3). Our results highlight the intricate relationship between structural chirality and optical activity in chiral crystals and caution against simplistic interpretations based solely on space group symmetry.

### v2

@@ Abstract
Natural optical activity (NOA), the ability of a material to rotate the plane of polarized light, has traditionally been associated with structural chirality. However, this relationship has often been oversimplified, leading to conceptual misunderstandings, particularly when attempts are made to directly correlate structural handedness with optical rotatory power. In reality, the relationship between chirality and NOA is more nuanced: optical activity can arise in both chiral and achiral crystal structures, and the sign of the rotation cannot necessarily be inferred from the handedness of the space group. In this work, we conduct a first-principles investigation of natural optical activity in SiO2 and AlPO4 crystals, focusing on their enantiomorphic structural phase transition from high-symmetry hexagonal ( or ) to low-symmetry trigonal ( or ) space groups. This transition, driven by the condensation of a zone-center phonon mode, reverses the screw axis type given by the space group symbol while leaving the sign of the optical activity unchanged. By following the evolution of the structure and the optical response along the transition pathway, we clarify the microscopic origin of this behavior. We demonstrate that the sense of optical rotation is determined not by the nominal handedness of the screw axis given in the space group symbol, but by the atomic-scale handedness of the most polarizable atoms of the structure.
@@ I Introduction
While it is a well-known fact that NOA is frequently linked with structural chirality, it is currently understood that chirality is a sufficient, albeit not necessary, condition for NOA to occur, given that achiral materials can also display NOA characteristics. [17, 18, 13]. This historical association led to the widespread belief that the handedness of a material’s space group directly determines the handedness of its optical rotation [19, 20]. Although this correlation may hold in many cases, it is more nuanced: the direction of optical rotation does not necessarily follow the handedness given in the space group symbol. [21, 22]
Measurements of optical activity are a common practice for organic chiral molecules (e.g., sugars or tartaric acid) and, although comparatively weaker, remain measurable in inorganic crystals such as low-quartz and berlinite [21]. Quartz-crystal (SiO2) adopts a trigonal structure (enantiomorphic space groups, 21 (152) or 21 (154)) at room temperature (“-quartz”) and transforms sharply to hexagonal “-quartz” (enantiomorphic space groups, (181) or (180) respectively) at 573 °C; both polymorphs are chiral despite being built from achiral SiO4 tetrahedra. Berlinite (AlPO4), the structural analogue of quartz, similarly undergoes an transition at approximately 586 °C, passing from trigonal () symmetry to hexagonal () without disrupting the network of corner-sharing tetrahedra [23, 24, 25]. In such transitions, the sense of the principal screw axis of the space group symbol reverses sign; however, the sense of the optical activity of the crystal remains unchanged [26, 22]. Initially, it might appear counterintuitive, but the direction of rotation for the chain composed of the most polarizable atoms (which is unchanged throughout the transition) actually determines the direction of optical activity, as shown by Glazer and Stadnicka. [21]. The arrangement of small- and large-channel helices endows quartz with a chiral bivector [27, 28]. This explains why the optical activity is preserved despite the reversal of the principal screw axis in these crystal structures.
@@ III.1 Structural transition of SiO2 and AlPO4
Recently, an algorithm has been proposed to identify chiral displacive transitions from achiral to chiral space groups [29], following the works of Fava et al. [38] for K3NiO2. Applying this workflow [29] to SiO2 and AlPO4, we find that a displacive phase transition from an achiral to a chiral space group is not allowed by pseudosymmetry. In other words, no achiral phase lies sufficiently close to the chiral phases of SiO2 or AlPO4 to permit a displacive transition.
@@
However, symmetry considerations predict that the high symmetry right-handed screw-axis structure can be linked to the low-symmetry 21 left-handed screw-axis structure [24, 25] through the irrep of the zone center [39, 30] (or similarly between the and the space groups). This distortion removes symmetry, leaving a screw as shown by the shaded hexagon and double triangles in Fig. 1. This means that those two space groups can form a high-symmetry/low-symmetry pairs for SiO2 or AlPO4.
@@
From our density functional theory (DFT) calculations, we confirm the presence of an unstable phonon mode at the point in the phase of SiO2, as referenced by [40]. By checking the corresponding phonon eigenvector symmetry, we found that it has the irrep, i.e. in agreement with the symmetry analysis which links the right-handed screw of the high-symmetry phase and the left-handed screw of the low-symmetry phase. Such a mode is neither Raman nor infrared active, i.e. it is silent as reported some time ago through hyper-Raman [41] or inelastic neutron scattering [42]. Hence, these results go in the direction of the displacive character of the transition [43], helping clarify previous discussions on its nature [44]. The calculated structural data of the converged structure computed with a -mesh of 666 can be found in Table 1. From our density functional perturbation theory (DFPT) calculations, we found that the unstable imaginary frequency of the mode has an amplitude of cm-1. The condensation and relaxation of the eigenvector gives an energy difference between the high and low temperature phases of meV per 9 atoms within the unit cell.
| Atom | x | y | z |
| Si (3d) | 0.0000 | 0.5000 | 0.3333 |
| O (6i) | 0.2080 | 0.4160 | 0.5 |
| Atom | x | y | z |
| Si (3a) | 0.0000 | 0.4689 | 0.6666 |
| O (6c) | 0.1442 | 0.7308 | 0.4515 |
@@
A symmetry adapted mode analysis using the isodistort software[45, 46] of our relaxed low-symmetry phase reveals that the distortion decomposes into and irreps where the amplitude of the mode is marginal (0.03 Å) with respect to the leading mode (1.00 Å) which drives the transition. The mode induces in-plane displacements of the Si atoms and both out-of-plane and in-plane displacements of the O atoms. The displacements are such that the SiO4 tetrahedra rotate as fairly rigid units (and do not distort significantly), as illustrated in Figure 1(c). As shown in Fig. 1(a-b), the in-plane projection of the displacements induces a coordinated counter-clockwise and clockwise movement between the neighboring helical centers. This movement breaks the -fold screw axis at the corners of the unit cells that transform into centers while preserving the -fold screw axis at the interior of the unit cell (see Table. 2).
@@
Moving now to the case of AlPO4 crystal and when studying its 22 phase, we can observe a similar crystal structure to that of the 22 of SiO2 [47, 48] with the exception that the -axis is essentially doubled due to the ordering of Al + P over the atom sites (as discussed below). The crystal structure is shown in Fig. 2. The calculated structural data of the converged AlPO4 22 structure can be found in Table 3. Similarly to the previous SiO2 case, we also obtain an unstable phonon mode at the point of the right-handed high-symmetry phase with an amplitude of cm-1 and the irrep [48]. This mode, hence, gives the link between the right-handed high-symmetry and the left-handed low-symmetry chiral phases, again similarly to the case of SiO2 but with doubled cell along the z direction. The condensation and relaxation of the eigenvector of this unstable mode gives, in this case, an energy gain of meV per 18-atom unit cell, which is comparable to SiO2 when normalized to the number of atoms. The relaxed structure we obtain is fully consistent with previous works [49, 47]. Similar conclusions to those drawn in the previous example can be reached by analyzing the distortion induced by the mode, which causes in-plane displacements of the Al and P atoms and both in-plane and out-of-plane displacements of the O atoms as in the case of the quartz crystal (see Table. 4). The symmetry adapted mode decomposition gives amplitudes of 0.08 Å for the irrep and 1.57 Å for the irrep.
@@ III.2 Evolution of the natural optical activity
The results are summarized in Fig. 3 and Table 5. The first interesting remark emerging from our calculations is the opposite sign of the NOA in the high-symmetry high-temperature phases of SiO2 and AlPO4, despite sharing the same enantiomorphic space group. Our first-principles calculations predict a positive rotatory power for the and phases of SiO2 for light traveling along its optical axis, in agreement with Ref. [11] (despite the handedness of the screw axis given in this space group symbol) and a negative rotatory power for light traveling along its optical axis for the and phases of AlPO4. The numerical values we obtained for SiO2 are in good agreement with the values encountered in the literature [50] and the change of rotatory sign between SiO2 and AlPO4 is also consistent with a previous report by Glazer and Stadnicka [21]. This finding underscores a crucial concept: the sign of a crystal’s natural optical activity links to its structural chirality, though it might not align with the helical handedness indicated by the space group symbol [21, 51].
@@
We can trace and explain the difference in rotatory power by examining the crystal structures of both materials. As mentioned in the previous section and now shown in Fig. 4, the atomic positions of AlPO4 in its setting are very similar to those of the SiO2 crystal in its setting. In fact, when we perform a computational experiment by substituting the Si atoms into the Al and P sites of the AlPO4 structure, and calculate the optical activity, we obtain a rotatory power with the opposite sign to that of SiO2 in its phase and that closely resembles that found for AlPO4 in its phase. Furthermore, upon relaxing the atomic positions in this modified structure, the system naturally evolves into the phase of SiO2 although initially a symmetry is detected by Abinit. Although the one-to-one alternation of Al and P atoms and the non-equivalence of the O atoms induces a right-handed screw axis of the material as discussed in Ref. [21, 47], Fig. 4(d) shows that a non-symmetric left-handed helix similar to that observed in 22 SiO2 is indeed present in the crystal. Although not perfectly symmetric, this chain formed by the most polarizable atoms in the crystal (oxygen sites in this particular case) dominates the optical activity of these phases explaining its rotatory power [21, 51]. Moreover, the observation that substituting Si atoms at the Al and P sites does not significantly alter the rotatory power suggests that these atomic species contribute little to the optical activity, which mainly arises from the more polarizable oxygen atoms.
The second interesting observation is that the condensation of the -distortion decreases the value of the rotatory power in these types of transitions where the screw-axis type is reverted. Indeed, as shown in Fig. 3 we observe a linear decrease of the rotatory power as we condense the distortion. This can be explained by looking at the distortions shown in Fig. 1 and Fig. 2 that tend to disrupt the helical arrangement of the O atoms. One might expect that further increasing the amplitude of the distortion could eventually revert the sign of the optical activity; however, our calculations show that, for instance, an amplitude 50% larger than the optimal value, recovers rotatory power values comparable to those of the undistorted high-symmetry phase. Moreover, this increase in distortion amplitude is accompanied by a significant increase in the Landau energy as shown in Fig. 3 (b), indicating that such configurations are physically inaccessible.
@@ IV Conclusion
In this work, we performed an ab-initio analysis of the structural transitions and associated changes in the natural optical activity between the high-symmetry () and low-symmetry () phases of SiO2 and AlPO4. Our first-principles calculations show that both materials undergo a displacive phase transition driven by the softening of a phonon mode at the zone center with symmetry. This mode induces a distortion that reverses the sign of the principal screw-axis of the crystal high-symmetry space group.
Despite this reversal, the sense of the natural optical activity of the materials is preserved throughout the transition. AlPO4 and SiO2 therefore provide a useful illustration of the principle that it is the handedness of the most polarizable atoms in a structure that determines the sign of its natural optical activity [21, 51]. Upon cooling through its phase transition, the screw-sense of the imperfectly symmetric helix of oxygen sites is not changed (although this is not obvious from the space group symmetries of the and phases where the helix handedness given by the space group symbol is reversed), preserving therefore the optical activity of the system.
@@
Furthermore, our analysis of the -transition pathway indicates a linear decrease in the magnitude of the rotatory power as the mode progresses from the high-symmetry to the low-symmetry phase for both the SiO2 and AlPO4 crystals. This trend suggests that the symmetry-lowering introduced by the distortion reduces the handedness of the structure (as shown in Figure 3). Our results highlight the intricate relationship between structural chirality and optical activity in chiral crystals and caution against simplistic interpretations based solely on space group symmetry.